Jimmy is making a rectangular garden. He wants it to be( x - 11 ) long and (x-1) yards wide. He needs a walkway 3 yards wide around the garden. If the total area of both garden and the walkway is 119 sq yards. solve for x.
Answer by ankor@dixie-net.com(22740)

Jimmy is making a rectangular garden. He wants it to be( x - 11 ) long and (x-1) yards wide.
He needs a walkway 3 yards wide around the garden.
If the total area of both garden and the walkway is 119 sq yards. solve for x.
:
From the information given, we know the overall dimensions will be
Length: x - 11 + 2(3) = (x - 5)
Width: x - 1 + 2(3) = (x + 5)
:
Overall area:
(x-5)*(x+5) = 119
FOIL
x^2 + 5x - 5x - 25 - 119 = 0
x^2 - 144 = 0
Factors to:
(x-12)(x+12) = 0
positive solution
x = 12
:
:
Check this by find the area with this solution
12-11+6 = 7
12-1 + 6 = 17
7 * 17 = 119